1.Draw a normal curve showing the 99% confidence interval

2.Appliace manufacturers are required to post a sticker on their products regarding the electricity economy of each appliance for sale. Explain how this sticker indicates estimation in general.

3.Compare and contrast the z-distribution and the t-distribution.

4.Using t-table, give the confidence coefficients for each of the following:

a. n = 12 with 95% confidence

b. n = 15 with 95% confidence

c. n = 21 with 99% confidence

d. n = 23 with 95% confidence

e. n = 25 with 99% confidence

5.Assuming that the samples come from normal distributions, find the margin of error E given the following:

a. n = 10 and X = 28 with s = 4.0, 90% confidence

b. n = 16 and X = 50 with s = 4.2, 95% confidence

c. n = 20 and X = 68.2 with s = 2.5, 90% confidence

d. n = 23 and X = 80.6 with s = 3.2, 95% confidence

e. n = 25 and X = 92.8 with s = 2.6, 99% confidence

6.Using the information in number 3, find the interval estimates of the population mean.

a.

b.

c.

d.

e.


1
Expert's answer
2021-03-31T13:20:03-0400

1.



2.

Dark green, it represents consumption of less than 25%.

With its intermediate green colour, it symbolizes consumption of less than 30%.

Its pale green colour represents consumption estimated as from 30% to 42%.

Yellow in colour, it offers average consumption between 42% and 55%.

Identified by the colour orange, it represents consumption of between 55% and 75%.

Bright orange in colour, it represents a consumption level of between 75% and 90%

Red in colour, it indicates that it has a consumption level of between 90% and 100%.


3.

The distribution curves are both​ mound-shaped and symmetric. ​ However, the​ t-test statistic curve is flatter than the​ z-statistic curve because the​ t-test is much more sensitive to the sample size. The standard normal or z-distribution assumes that you know the population standard deviation. The t-distribution is based on the sample standard deviation.


4.a) "t=1.78"

b) "t=1.75"

c) "t=2.52"

d) "t=1.71"

e) "t=2.49"


5.

"E=z_{\alpha}\frac{s}{\sqrt{n}}"

a.

"E=1.65\cdot\frac{4}{\sqrt{10}}=2.087"

b.

"E=1.96\cdot\frac{4.2}{\sqrt{16}}=2.058"

c.

"E=1.65\cdot\frac{2.5}{\sqrt{20}}=0.922"

d.

"E=1.96\cdot\frac{3.2}{\sqrt{23}}=1.308"

e.

"E=2.58\cdot\frac{2.6}{\sqrt{25}}=1.3416"


6.

"z_{\alpha/2}<\frac{X-\mu}{s/\sqrt{n}}<z_{1-\alpha/2}"

"(X-z_{\alpha/2}\cdot s/\sqrt{n})>\mu>(X-z_{1-\alpha/2}\cdot s/\sqrt{n})"

a.

"(28+2.575\cdot4/\sqrt{10})>\mu>(28-1.645\cdot4/\sqrt{10})"

"25.92<\mu<31.26"

b.

"(50-1.96\cdot4.2/\sqrt{16})<\mu<(50+1.96\cdot4.2/\sqrt{16})"

"47.94<\mu<52.06"

c.

"(68.2-1.645\cdot2.5/\sqrt{20})<\mu<(68.2+2.575\cdot2.5/\sqrt{20})"

"67.28<\mu<69.64"

d.

"(80.6-1.96\cdot3.2/\sqrt{23})<\mu<(80.6+1.96\cdot3.2/\sqrt{23})"

"79.29<\mu<81.91"

e.

"(92.8-2.575\cdot2.6/\sqrt{25})<\mu<(92.8+2.575\cdot2.6/\sqrt{25})"

"91.46<\mu<94.14"